OpenStudy (anonymous):

Add six and one-sixth plus two and two-thirds. nine and five-sixths eight and two-thirds nine and two-thirds eight and five-sixths

OpenStudy (anonymous):

help

OpenStudy (anonymous):

\[1 \frac{ 1 }{ 6 } + 2 \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

yep that's it

OpenStudy (anonymous):

@Hero can u help, now that i have a problem

OpenStudy (anonymous):

rose, do u understand?

OpenStudy (anonymous):

the question?

OpenStudy (anonymous):

\[\frac{ 7 }{ 6 }+\frac{ 8 }{ 3 }\] \[\frac{ 8 }{ 3 } \times \frac{ 2 }{ 2 }\] \[\frac{ 7 }{ 6 } + \frac{ 16 }{ 6 }\]

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

Add across on the top and the bottom stays the same.

OpenStudy (anonymous):

DUDE!!! im in 6th grade plz explain more simplifiedly

OpenStudy (anonymous):

please type faster

OpenStudy (anonymous):

So first I'm not a dude Second do not tell me what to do. You're lucky I decided to even help you after you demanded me to type faster. You have to make the denominators(bottom) the same. I did that by multiplying the numerator and denominator by 2 That's how I got 16 over 6. So now that the denominators are the same you add the numerators which would be 7 + 16 = 23 so we now have 23 over 6 6 goes into 23 3 times with a remainder of 5 so the answer is 3 5 over 6

OpenStudy (anonymous):

okay sorry, i just cant be here for so long

OpenStudy (anonymous):

thank you so much and have a great day

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

that is not an answer,

OpenStudy (anonymous):

And I know why lol The answer is actually 8 and 5 over 6

OpenStudy (anonymous):

oh... ok

OpenStudy (anonymous):

thanx